A company is buying machinery for their manufacturing floor that is worth $455,000. The machinery has a service life of 7 years. Develop a depreciation table for this asset according to MACRS.

Please show all work including formulas used. Thanks!
Answer #1

Depreciation Depreciation year Amount Rate 1 455000 14.29% | 2 455000 24.49% | 3 | 455000 17.49% 4 455000 12.49% 5 | 455000 8

We can get the MACRS rate from the table and then multiply the rate with the amount.

For example in first year depreciation will be equal to

455,000 × 0.1429 = $ 65,019.50

Like wise we can calculate for all the years.

BV at the end of year

First year, 455,000 - 65,019.50 = $ 389,980.50

Second year, $ 389,980.50 - 111,429.50 = $ 278,551

Third year, $ 278,551 - 79,579.50 = 198,971.50

Likewise we can calculate the EOY for all the years.

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.